# ELFuzz: Efficient Input Generation via LLM-driven Synthesis Over Fuzzer Space

### v3

@@ Abstract
Generation-based fuzzing produces appropriate testing cases according to specifications of input grammars and semantic constraints to test systems and software. However, these specifications require significant manual efforts to construct. This paper proposes a new approach, ELFuzz (Evolution Through Large Language Models for Fuzzing), that automatically synthesizes generation-based fuzzers tailored to a system under test (SUT) via LLM-driven synthesis over fuzzer space. At a high level, it starts with minimal seed fuzzers and propels the synthesis by fully automated LLM-driven evolution with coverage guidance. Compared to previous techniques, ELFuzz can 1) seamlessly scale to SUTs of real-world sizes (up to 1,791,104 lines of code in our evaluation) and 2) synthesize efficient fuzzers that catch interesting grammatical structures and semantic constraints in a human-understandable way. Our evaluation shows that ELFuzz achieves up to 418.5% more coverage and triggers up to 216.7% more artificially injected bugs, compared to the state-of-the-art. We also used ELFuzz to conduct a real-world fuzzing campaign on the newest version of cvc5 for 14 days, and encouragingly, it found five 0-day bugs (three are exploitable). Moreover, our ablation study shows that fuzzer space, the key component of ELFuzz, contributes the most (up to 62.5%) to the effectiveness of ELFuzz. Further analysis of the fuzzers synthesized by ELFuzz confirms that they correctly express the grammatical structures and semantic constraints of a SUT. The results present a promising potential of ELFuzz for more automated, efficient, and extensible input generation for fuzzing.
@@ 3     $\mathrm{cov}_{m}\leftarrow$$\textsc{ApproxCov}_{SUT,T}$($m$)
[…]
@@ 2   $\mathrm{attempts}\leftarrow\emptyset$
4     $\mathrm{attempt}_{i}\leftarrow$$\textsc{GreedyMax}_{M,N}$($\mathrm{covSets}$)
@@ 5     $\mathrm{attempts}\leftarrow\mathrm{attempts}\cup\{\mathrm{attempt}_{i}\}$
12       $\mathrm{attempt}\leftarrow$$\textsc{RandPick}_{N}$($\mathrm{covSets}$)
@@ 6.4.1 RQ1: Effectiveness in fuzzing
Figure 7 presents the edge coverage of inputs generated by the four fuzzers within ten minutes. For every SUT, ELFuzz achieves the best performance, with up to 434.8% more covered edges (on the cvc5 benchmark). There are no results for ISLearn on the re2 benchmarks because ISLearn failed to synthesize any semantic constraints for it (possibly because its predefined templates do not apply to the SUT). This shows the limited applicability of ISLearn.
@@
Figure 8 presents the coverage trends of AFL++ on seeds generated by each fuzzer to compare. The solid lines are the averaged values over the 10 repetitions, and the narrow shadows show the standard deviations. Though the parts other than the seeds are identical in the fuzzing processes, the high coverage of seeds generated by ELFuzz brings significant advantages to the later fuzzing processes. On libxml2, CPython, SQLite, cvc5, and librsvg, AFL++ instances that use other seeds can never catch up to the instance that uses the ELFuzz seeds, and the gaps stay significant (up to 436.4% when comparing ELFuzz with the second-best results) at the end of the experiments. On jsoncpp and re2, other techniques gradually reach similar coverage as time elapses. This is because the JSON format and regular expressions both have relatively simple grammars. The grammatical tokens of them are mostly single characters (e.g., “{” and “[” in JSON, and “?”, “*” in regular expressions) and do not require complex rules (e.g., matching tags in XML) to be valid. Mutation-based fuzzing can work well by itself, whatever the seeds are; therefore, the gulf brought by the seeds is gradually bridged when given sufficient time. However, seeds generated by ELFuzz reduce the time to reach high coverage.
@@ 6.4.2 RQ2: Effectiveness in bug finding
Figure 9 presents the number of bugs triggered by AFL++ within 24 hours using different seeds. The solid lines are the average values, and the shadows show the standard deviations. The results demonstrate that the seeds generated by ELFuzz can not only find more bugs but also find the bugs faster. Table 6 presents the time (averaged over the 10 repetitions) required by the seeds generated by each fuzzer to reach 25%, 50%, and 75% of the number of all bugs triggered by the seeds generated by the ELFuzz fuzzers as quantitative metrics to demonstrate the different speeds.
@@ 7 Limitations and Future Work
Currently, ELFuzz acts as the input generator for initial seeds. Thus, it mainly boosts the early stage of the fuzzing cycle. However, maintaining corpus diversity is crucial during the entire fuzzing process. Extending the use of the ELFuzz fuzzers to the entire fuzzing cycle could benefit fuzzing processes more, but it faces a main difficulty caused by the complex dynamics of fuzzing. As the fuzzer touches different parts of a SUT, the desired properties of the corpus keep changing. For example, the importance of grammatical structures may decrease after the fuzzer proceeds to the assembly optimization passes when testing a compiler. Moreover, rarely used features initially ignored by ELFuzz may be discovered, and covering them would require adaptation of the ELFuzz fuzzer to cover them. A possible solution is to continuously evolve the input generators throughout the entire fuzzing cycle to keep pace with such changing needs. The continuous evolution and input generation can also make the time spent on fuzzer synthesis before the fuzzing campaigns more cost-effective. Future research may investigate how to leverage runtime feedback to facilitate continuous evolution, enabling adaptive input generation and the incorporation of newly discovered features of the SUTs while maintaining reasonable resource costs (such as GPU hours).
@@ Appendix B Failure Case Analysis
[…]
@@
[…]
@@ Appendix C The Real-world Bugs in cvc5
    Bug 5: Crash in nullable checks without arguments. This bug triggers a NULL-pointer dereference when calling nullable checks without arguments.

### v4

@@ Abstract
Generation-based fuzzing produces appropriate testing cases according to specifications of input grammars and semantic constraints to test systems and software. However, these specifications require significant manual effort to construct. This paper proposes a new approach, ELFuzz (Evolution Through Large Language Models for Fuzzing), that automatically synthesizes generation-based fuzzers tailored to a system under test (SUT) via LLM-driven synthesis over fuzzer space. At a high level, it starts with minimal seed fuzzers and propels the synthesis by fully automated LLM-driven evolution with coverage guidance. Compared to previous techniques, ELFuzz can 1) seamlessly scale to SUTs of real-world sizes (up to 1,791,104 lines of code in our evaluation) and 2) synthesize efficient fuzzers that catch interesting grammatical structures and semantic constraints in a human-understandable way. Our evaluation shows that ELFuzz achieves up to 434.8% more coverage over the second best and triggers up to 216.7% more artificially injected bugs, compared to the state-of-the-art. We also used ELFuzz to conduct a real-world fuzzing campaign on the newest version of cvc5 for 14 days, and encouragingly, it found five 0-day bugs (three are exploitable). Moreover, our ablation study shows that fuzzer space, the key component of ELFuzz, contributes the most (up to 62.5%) to the effectiveness of ELFuzz. Further analysis of the fuzzers synthesized by ELFuzz confirms that they correctly express the grammatical structures and semantic constraints of a SUT. The results present a promising potential of ELFuzz for more automated, efficient, and extensible input generation for fuzzing.
@@ 4.2 Fuzzer Space Exploration
Take the seed fuzzers $\mathrm{F}_{\mathrm{B}}$, $\mathrm{F}_{\mathrm{D}}$ and the mutants $\mathrm{F}_{\mathrm{H}}$$\mathrm{F}_{\mathrm{K}}$ in the previous subsection as an example. After running for $T$ time, $\mathrm{F}_{\mathrm{J}}$ triggers type errors and is thus discarded. The cover sets of the other fuzzers are: 1) lines 15 and 13 for $\mathrm{F}_{\mathrm{B}}$ producing (*"(" , (*"((" and (*"" , 2) lines 13 and 1113 for $\mathrm{F}_{\mathrm{D}}$ producing (*"*" , (*"**" and (*"" , 3) lines 16 and 1113 for $\mathrm{F}_{\mathrm{G}}$ producing (*"(" , (*"*" and (*"" , 4) lines 13 and 13 for $\mathrm{F}_{\mathrm{I}}$ producing (*"" , and 5) lines 17 and 910 for $\mathrm{F}_{\mathrm{K}}$ producing (*"(" , (*")" and (*")(" .
@@ 2   $\mathrm{attempts}\leftarrow\emptyset$
[…]
@@ 5     $\mathrm{attempts}\leftarrow\mathrm{attempts}\cup\{\mathrm{attempt}_{i}\}$
[…]
@@ 6.4.1 RQ1: Effectiveness in fuzzing
Figure 7 presents the edge coverage of inputs generated by the four fuzzers within ten minutes. For every SUT, ELFuzz achieves the best performance, with up to 434.8% more covered edges over the second best (on cvc5). There are no results for ISLearn on the re2 benchmarks because ISLearn failed to synthesize any semantic constraints for it (possibly because its predefined templates do not apply to the SUT). This shows the limited applicability of ISLearn.
@@
Figure 8 presents the coverage trends of AFL++ on seeds generated by each fuzzer to compare. The solid lines are the averaged values over the 10 repetitions, and the narrow shadows show the standard deviations. Though the parts other than the seeds are identical in the fuzzing processes, the high coverage of seeds generated by ELFuzz brings significant advantages to the later fuzzing processes. On libxml2, CPython, SQLite, cvc5, and librsvg, AFL++ instances that use other seeds can never catch up to the instance that uses the ELFuzz seeds, and the gaps stay significant (up to 436.4% compared to the second-best results) at the end of the experiments. On jsoncpp and re2, other techniques gradually reach similar coverage as time elapses. This is because the JSON format and regular expressions both have relatively simple grammars. The grammatical tokens of them are mostly single characters (e.g., “{” and “[” in JSON, and “?”, “*” in regular expressions) and do not require complex rules (e.g., matching tags in XML) to be valid. Mutation-based fuzzing can work well by itself, whatever the seeds are; therefore, the gulf brought by the seeds is gradually bridged when given sufficient time. However, seeds generated by ELFuzz reduce the time to reach high coverage.
@@ 6.4.2 RQ2: Effectiveness in bug finding
Figure 9 presents the number of bugs triggered by AFL++ within 24 hours using different seeds. The solid lines are the average values, and the shadows show the standard deviations. The results demonstrate that the seeds generated by ELFuzz can not only find more bugs but also find the bugs faster. At the end, ELFuzz triggers up to 216.7% more bugs (on SQLite over GLADE). Table 6 presents the time (averaged over the 10 repetitions) required by the seeds generated by each fuzzer to reach 25%, 50%, and 75% of the number of all bugs triggered by the seeds generated by the ELFuzz fuzzers as quantitative metrics to demonstrate the different speeds.
@@ 7 Limitations and Future Work
Currently, ELFuzz acts as the input generator for initial seeds. Thus, it mainly boosts the early stage of the fuzzing cycle. However, maintaining corpus diversity is crucial during the entire fuzzing process. Extending the use of the ELFuzz fuzzers to the entire fuzzing cycle could benefit fuzzing processes more, but it faces a main difficulty caused by the complex dynamics of fuzzing. As the fuzzer touches different parts of a SUT, the desired properties of the corpus keep changing. For example, the importance of grammatical structures may decrease after the fuzzer proceeds to the assembly optimization passes when testing a compiler. Moreover, rarely used features initially ignored may be discovered, and covering them would require adaptation of the ELFuzz fuzzer to cover them. A possible solution is to continuously evolve the input generators throughout the entire fuzzing cycle to keep pace with such changing needs. The continuous evolution and input generation can also make the time spent on fuzzer synthesis before the fuzzing campaigns more cost-effective. Future research may investigate how to leverage runtime feedback to facilitate continuous evolution, enabling adaptive input generation and the incorporation of newly discovered features of the SUTs while maintaining reasonable resource costs (such as GPU hours).
@@ Appendix B Failure Case Analysis
[2.5pt/5pt] CPython 5063, 5094, 5695 These three bugs are in code that handles keyword arguments. The ELFuzz fuzzers and the ANTLR4 grammars can both generate test cases containing this feature.
@@
[2.5pt/5pt] SQLite 469 The bug is in the code that handles the round function. Neither ELFuzz fuzzers nor the ANTLR4 grammars specify such cases explicitly, while the mutation process may hit it by chance.
@@ Appendix C The Real-world Bugs in cvc5
    Bug 5: Crash in Tnullable checks without arguments. This bug triggers a NULL-pointer dereference when calling Tnullable checks without arguments.
@@ Appendix E Extending ELFuzz with Zest
USENIX Security ’25 Artifact Appendix ELFuzz: Efficient Input Generation via LLM-driven Synthesis Over Fuzzer Space
Appendix A Artifact Appendix
This artifact appendix is a self-contained document that describes the roadmap for the evaluation of the artifacts of the paper “ELFuzz: Efficient Input Generation via LLM-driven Synthesis Over Fuzzer Space.” It includes the hardware, software, and configuration requirements for the evaluation, as well as the process for reproducing the results and claims presented in the paper.
The artifacts have acquired the “artifact available” and “artifact functional” badges after the artifact evaluation. The full replication experiments for the “artifact reproduced” badge were not finished within the one-month time span of the artifact evaluation process due to the requirements of a significant amount of time and substantial computational resources. However, our testing shows that the artifacts function well to reproduce all the results in the paper with enough time and computational resources provided. Furthermore, we will continue to improve the artifacts for better usability and bug fixes after the artifact evaluation.
A.1 Abstract
The artifacts contain the implementation of ELFuzz, the source code of the benchmarks and baselines compared with ELFuzz in the evaluation, the evaluation results, and the data required to reproduce the evaluation. It also includes a Docker image that preserves the environment and setups of the evaluation to facilitate one-touch replication.
A.2 Description & Requirements
This section outlines the hardware and software requirements necessary to recreate the experiment environment and setups described in the paper, as well as the steps to acquire the benchmarks and baselines used in the evaluation.
A.2.1 Security, privacy, and ethical concerns
The artifact does not perform any operations that may compromise the security or privacy of the evaluators who execute it.
A.2.2 How to access
The replication package, containing the code and data, is published on Zenodo. The download link of the version at the time of artifact evaluation is https://doi.org/10.5281/zenodo.16741080. Updated versions will be hosted at https://doi.org/10.5281/zenodo.15833146. The source code of ELFuzz and the replication package is developed in the GitHub repository at https://github.com/OSUSecLab/elfuzz.
A.2.3 Hardware dependencies
The experiments require a GPU with 26 GiB VRAM and CUDA support to run a local CodeLlama-13B model. The intermediate data and results produced or downloaded by the experiments will occupy approximately 100 GiB of disk space. In our evaluation, we use 30 processes to accelerate the experiments. It is recommended that the evaluators’ machines be equipped with CPUs with at least this number of cores. Fewer CPU cores do not affect the results of the experiments, but a longer time to finish the experiments should be expected. The specific GPU and CPUs used in the original evaluation are one NVIDIA H100 Tensor Core GPU and two AMD EPYC 7251 CPUs.
A.2.4 Software dependencies
The experiments do not depend on a specific operating system. A Docker installation that is compatible with version 24.0.7 is required. The host machine should also have the CUDA toolkit installed.
A.2.5 Benchmarks
Benchmarks used in the evaluation will be automatically cloned from their official Git repositories or downloaded from the official sites during the experiments. Dockerfiles from the Fuzzbench project [^1] https://github.com/google/fuzzbench and the OSS-Fuzz project [^2] https://github.com/google/oss-fuzz are modified for replicable building environments for the six benchmarks other than cvc5. The Dockerfile for cvc5 is implemented from scratch. All Dockerfiles are included in the source code tarball of the artifacts.
A.3 Setup
This section describes how to set up and configure the environment to replicate the evaluation. It also includes instructions on the functionality validation of the replication package.
A.3.1 Installation
First, follow the official instructions [^3] https://docs.docker.com/engine/install/ to install Docker. Then, set up the core pattern in the host machine as required by AFL++ later. Hereafter, “\$” indicates user inputs, and “>” indicates program outputs.
    \$ echo core > /proc/sys/kernel/core_pattern
Now, import the Docker image and launch the container:
    \$ zstd -d elfuzz_docker_<timetag>.tar.zst
    \$ docker load --input elfuzz_docker_<timetag>.tar
    \$ mkdir -p /tmp/host
    \$ docker run --storage-opt size=100G \
                 --cpus 30 \
                 --add-host=host.docker.internal:host-gateway \
                 -v /tmp/host:/tmp/host \
                 -v /var/run/docker.sock:/var/run/docker.sock \
                 --name elfuzz \
                 -it ghcr.io/osuseclab/elfuzz:25.07.2
The commands should lead you into the container where the experiments will happen.
A.3.2 Initilizing the environment
After entering the Docker container, run the following commands to enable sibling containers.
    \$ sudo chown -R appuser:appuser /tmp/host/
    \$ elfuzz setup
    \$ exit
The two commands will exit the container. Now, restart the container for the settings to take effect.
    \$ start -ai elfuzz
Then, the following command will download all the data files from Zenodo and place them in the correct locations.
    \$ elfuzz download
You also need to configure your Hugging Face token:
    \$ elfuzz config --set tgi.huggingface_token <YOUR_TOKEN>
A.3.3 Functionality Validation
After initializing the environment, you can run mini-versions of the experiments presented in the paper to validate that all of them function well. These commands adopt the same settings as the original experiments, but decrease the parameters such as time limits and the number of evolution iterations for quick validation. Each command can take 5 minutes to 2 hours to complete. §A.4 will give detailed explanations of these commands.
    \$ elfuzz synth -T fuzzer.elfuzz \
                   --use-small-model \
                   --evolution-iterations 3 \
                   jsoncpp
    \$ elfuzz synth -T grammar.glade re2
    \$ elfuzz synth -T semantics.islearn cvc5
    \$ elfuzz produce --time 10 -T glade jsoncpp
    \$ elfuzz minimize -T glade jsoncpp
    \$ elfuzz run rq1.seed_cov -T glade jsoncpp
    \$ elfuzz run rq1.afl --fuzzers glade \
                         --repeat 1 \
                         --time 60 \
                         jsoncpp
    \$ elfuzz run rq2.afl --fuzzers glade \
                         --repeat 1 \
                         --time 60 \
                         jsoncpp
    \$ elfuzz run rq2.triage
    \$ elfuzz run rq2.real_world --time 60
    \$ elfuzz run rq3
These commands should output messages like “ELFuzz fuzzers successfully synthesized” without being interrupted by errors.
A.4 Reproducibility Validation
This section outlines the major claims presented in the evaluation part of the paper (§A.4.1). It describes how to conduct the experiments to fully reproduce the results presented in the paper, thereby supporting these claims (§A.4.2). However, these experiments require a significant amount of time and substantial computational resources. Thus, another option is to run the smaller-scale experiments (§A.4.2), which share the same settings as the full-scale experiments but are conducted for a shorter time or with fewer repetitions. Results produced by the small-scale experiments should also support our claims, although not 100% replications of those in our original paper. We will explain how to conduct them and draw the same claims from the results.
A.4.1 Major claims
The major claims of the paper are as follows:
  • $\bullet$
    C1. Fuzzers synthesized by ELFuzz significantly outperform the state-of-the-art generation-based fuzzers, respecting the coverage of the produced test cases and the promotion that the test cases bring to later mutation-based fuzzing processes. This is proven by the input generation experiments (E1) and the mutation-based fuzzing experiment (E2) in RQ1.
  • $\bullet$
    C2. Fuzzers synthesized by ELFuzz significantly outperform the state-of-the-art generation-based fuzzers when being used to find artificially injected bugs, and they can find previously unknown bugs in real-world SUTs. This is proved by the bug-finding experiments on bug-injected benchmarks (E3) and the real-world bug-finding experiment on cvc5 (E4).
  • $\bullet$
    C3. Fuzzer space contributes the most to the performance of ELFuzz among all the components. This is proved by the input generation experiments (E5) and the LLM-driven evolution processes (E6).
  • $\bullet$
    C4. Fuzzers synthesized by ELFuzz are interpretable and extensible. This is proved by case studies on the fuzzer code (E7) and manual adaptation of Zest, another input-generation technique, onto the fuzzers (E8).
A.4.2 Full-scale experiments
Below are instructions on conducting full-scale experiments to replicate the results in the paper. It also lists the expected outputs and how they support our claims. Note that if you have validated the functionality of the replication package before the experiments, please reset the Docker container (remove, re-launch, and re-initialize it) to avoid data pollution caused by the results produced by the functionality validation process.
  • $\bullet$
    E1, E5, and E6. [1 human-hour + 10 compute-day + 50GiB disk] Inside the Docker container, use the following command to synthesize the fuzzers using ELFuzz or its variants (E6), or mine grammars and semantic constraints using the baselines:
        \$ elfuzz synth -T <baseline> <benchmark>
    Please refer to the help message (--help) on what values <baseline> and benchmark can take. The following commands conduct the input generation experiments (E1 and E5):
        \$ elfuzz produce -T <baseline> <benchmark>
        \$ elfuzz minimize -T <baseline> <benchmark>
        \$ elfuzz run rq1.seed_cov -T <baseline> <benchmark>
        \$ elfuzz run rq3
    The coverage of the generated seed test cases will be recorded in an Excel sheet, which can be viewed via
        \$ pyexcel view /elfuzz/analysis/rq1/results/seed_cov.xlsx
    The coverage of seed test cases generated by ELFuzz is expected to be significantly higher than that generated by other methods (E1 results for C1). Similarly, you can view the coverage of seed test cases generated by ELFuzz variants via
        \$ pyexcel view /elfuzz/analysis/rq3/rq3_ablation.xlsx
    The coverage of seed test cases generated by ELFuzz-noSP, ELFuzz-noCP, and ELFuzz-noIN will show slight decreases, while that generated by ELFuzz-noFS will show significant decreases (E5 results for C3). The coverage trends of candidate fuzzers of the four variants during evolution can be viewed via
        \$ pyexcel view /elfuzz/analysis/rq3/rq3_evolve_cov.xlsx
    The values will display curves similar to those shown in Figure 12 of the paper if being drawn (E6 results for C3).
  • $\bullet$
    E2 and E3. [1 human-hour + 20 compute-day + 5GiB disk] The following command runs the mutation-based fuzzing experiments in RQ1 (E2) and collects and analyzes the results:
        \$ elfuzz run rq1.afl --fuzzers <baseline_list> \
                             --repeat 10 <benchmark_list>
    The following commands will run the mutation-based fuzzing campaign in RQ2 (E3) and collect and analyze the results:
        \$ elfuzz run rq2.afl --fuzzers <baseline_list> \
                             --repeat 10 <benchmark_list>
        \$ elfuzz run rq2.triage
    You can view the averaged coverage trends during AFL++ fuzzing campaigns for RQ1 and trends of triggered bugs via
        \$ pyexcel view /elfuzz/rq1/results/rq1_sum.xlsx
        \$ pyexcel view /elfuzz/rq2/results/rq2_count_bugs.xlsx
    The data will show curves similar to those in Figures 8 and 9, i.e., the curves of ELFuzz will be higher than that of others (E2 for C1 and E3 for C2).
  • $\bullet$
    E4. [8 human-hour + 14 compute-day] The following command runs the real-world bug-finding experiment:
        \$ elfuzz run rq2.real_world
    The triage and analysis of the results can only be done manually. The data tarball contains the bug-triggering test cases we found, though, in rq2/results/cvc5_bugs. You can check whether they actually trigger the bugs of the corresponding versions of cvc5 (E4 for C2).
  • $\bullet$
    E7 and E8. These two empirical case studies were done manually. The README file in the Docker image contains suggestions on how to replicate them (E7 and E8 for C4).
  • $\bullet$
    Reproducing the figures and tables. Use the following command to reproduce the figures and plots presented in the paper:
        \$ elfuzz plot --all
    Note that we cannot include proprietary fonts in the replication package. The appearance of the final rendered figures may be different from that in the original paper.
Please refer to the README files in the Zenodo repository, source code tarball, and Docker image tarball for more details.
A.4.3 Small-scale experiments
The full-scale experiments require a significant amount of time and substantial computational resources. Instead, you can choose to run some experiments at smaller scales, and the results can also support our claims. Below are instructions.
  • $\bullet$
    E1 and E5. [1 human-hour + 1 compute-day + 25GiB disk] Unfortunately, decreasing the number of evolution iterations will significantly affect the performance of the synthesized fuzzers, which is expected. Thus, we cannot run the small-scale versions of the synthesis processes and E6. Therefore, we will utilize the original fuzzers included in the replication package. For E1 and E5, you can use the --time 60 option for the produce subcommand to run the seed test case generation for 1 minute instead of 10. Even this short time limit should already demonstrate the advantage of ELFuzz (C1) and the huge impact of excluding the fuzzer space from ELFuzz (C3).
  • $\bullet$
    E2 and E3. [1 human-hour + 5 compute-day + 5GiB disk] Note that using seed test cases generated in a time shorter than 10 min can weaken the advantage of ELFuzz in AFL++ fuzzing campaigns, which is expected. Please reset the Docker container if you have run the above small-scale experiments before starting these two experiments, to ensure that you use the original test cases (generated in 10 minutes) included in the replication package. Then, you can change the value of the --repeat option from 10 to 3 to decrease the number of repetitions. Use the --time 3600 option to run the AFL++ fuzzing campaigns for 1 hour instead of 24 hours. ELFuzz should be able to show advantages within the first hour of fuzzing.
A.5 Version
Based on the LaTeX template for Artifact Evaluation V20231005. Submission, reviewing and badging methodology followed for the evaluation of this artifact can be found at https://secartifacts.github.io/usenixsec2025/.